OpenStudy (anonymous):

omg i forgot, what is the special rule to calculate 2 cos value again? like for sin its: sin(x) = sin(pi-x) whats cos? cos(x) = ?

OpenStudy (mathmate):

cos(x)=cos(-x)

OpenStudy (anonymous):

@mathmate love yo a$$ ;)

OpenStudy (mathmate):

:)

OpenStudy (anonymous):

@mathmate and what about Tan?

OpenStudy (mathmate):

tan(x)=tan(x+Pi)=tan(x-Pi)